OpenStudy (anonymous):

f(2n)= 2f(n) for all in integers n f(4) =4 IF f is a function defined for all positive integers n, and f satisfies the two conditions above, which of the following could be the definition of f?

OpenStudy (anonymous):

let n=2 f(4)=2f(2) 4=2f(2) f(2)=2 let n=1 f(2)=2f(1) 2=2f(1) f(1)=1 ............................... it seems\[f(n)=2^{n-1}\]is an answer...we prove by induction \[f(1)=2^{1-1}=1\]\[f(n+1)=2f(n)=2\times 2^{n-1}=2^n\]

OpenStudy (anonymous):

so \[f(n)=2^{n-1}\]must be one of ur options